OpenStudy (anonymous):

Find the 10th term of the following sequence: -5, -10, -20 . . .

OpenStudy (anonymous):

Find the 10th term of the following sequence: -5, -10, -20 . . . 2 • (-5)9 2 • (-5)10 -5 • 29 -5 • 210

OpenStudy (anonymous):

the n'th term of a geometric sequence is a*r^(n-1), where a is the first term and r is the common ratio so 10th term is -5*2^(9) = -5*64*8 = -320*8 = -2560

OpenStudy (anonymous):

SO wait which one is it out of the list

OpenStudy (anonymous):

it's none of them...it's -5*512